Shift report 20030322 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Mar 22 2003 - 15:47:21 EST

  • Next message: Apache: "Shift report 20030322 16:00-24:00"
    ------------------------------------------------------------
    Date: Sat Mar 22 15:47:20 2003, Shift: 08:00-16:00
    Supervisor: SL, with: 
    Log book#14 pages 72 - 72
    
    Shift Summary:
    ------------------------------------------------------------
    1.Run 8604 ongoing at start of shift. Store ended after votes from brahms and phenix. No beam rest of the shift due to cryo-problems. 
    2.Run 8604 completed 40k tr6 events with fs at 12deg D1=D2=1/8; D3=D4=0. Runs 8605 through 8608 are pedestal and monitor runs.
    3.No new detector problems.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sat Mar 22 2003 - 15:48:19 EST